The height of a projectile fired upward is given by the formula
s = v0t − 16t2,
where s is the height, v0 is the initial velocity, and t is the time. Find the time for a projectile to return to Earth if it has an initial velocity of 248 ft/s.

Answer by ewatrrr(24785)


Hi
The height of a projectile fired upward is given by the formula:
s = v0t − 16t^2 |Note the use of " ^ "
Find the time for a projectile to return to Earth if it has an initial velocity of 248 ft/s.
s = 0 when the projectile returns to earth.
0 = 248t - 16t^2
16t^2 = 248t
t = 248/16 = 15.5s
Wish You the Best in your Studies.